# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v3

@@ IV Complexity analysis
Under Assumptions A1-A2 and A4-A10, along with conditions of Corollary 2 and Lemma 4, there exists some $1\leq t\leq{\mathcal{O}}(\epsilon^{-3/2})$ and ${\hat{{\boldsymbol{\lambda}}}}_{t}\in{\mathbb{R}}^{I}_{+}$, ${\hat{{\boldsymbol{\nu}}}}_{t}\in{\mathbb{R}}^{J}_{+}$ such that the point $({{\hat{{\mathbf{x}}}}_{t}},{\hat{{\boldsymbol{\lambda}}}}_{t},{\hat{{\boldsymbol{\nu}}}}_{t})$ is $\epsilon$-KKT optimal.
Under the adaptive step-size assumption, the SFO complexity bound in Theorem 2 matches that of STORM [20] and nearly matches Pstorm [33], STORM+ [30] achieving a complexity of $\mathcal{O}(\epsilon^{-3/2})$ for the unconstrained case. Notably, while STORM and Pstorm were proposed for unconstrained and convex-constrained problems, respectively, CoSTA extends this to non-convex constrained framework with a comparable SFO complexity. Furthermore, under the non-adaptive step-size setting, CoSTA attains the optimal convergence rate established by [14]. A key distinction of CoSTA, however, lies in its ability to achieve a near-optimal rate (i.e., ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$) under adaptive step sizes, a feature that distinguishes it from [14], which employs non-adaptive step sizes. It is remarked that the non-asymptotic results in Theorem 2 and Corollary 3 can also be used to infer the asymptotic convergence of the corresponding quantities.
@@ V-A Sparse Binary Classification
The problem of finding a sparse binary classifier can be written in the form of ($\mathcal{P^{\prime}}$) with the objective being the logistic loss function, i.e., $f({\mathbf{x}},\{{\mathbf{a}},b\})=\log(1+e^{-b{\mathbf{a}}^{\mathsf{T}}{\mathbf{x}}})$ and the constraint $g({\mathbf{x}})$ taking form of a smoothed sparsity inducing function, such as the $\ell_{1}$ norm or the minimax concave function [9]. The random variable here signifies the independent observations $\{{\mathbf{a}},b\}$. For our case we will take $\bar{g}$ to be minmax concave penalty (MCP) [9], given by:
$\bar{g}({\mathbf{x}})$ $=\lambda\|{\mathbf{x}}\|_{1}-\sum_{k=1}^{n}\bar{h}_{\lambda,\theta}([{\mathbf{x}}]_{k})$ (65)
where $[{\mathbf{x}}]_{k}$ denotes the $k$-th entry of ${\mathbf{x}}$ and $\bar{h}_{\lambda,\theta}$ is defined as:
$\bar{h}_{\lambda,\theta}(x)=\begin{cases}\frac{x^{2}}{2\theta},&|x|\leq\theta\lambda\\ \lambda|x|-\frac{\theta\lambda^{2}}{2},&|x|>\theta\lambda.\\ \end{cases}$ (66)
It can be noted that, even though $\bar{g}$ is non-smooth, a smoothed version of the MCP can be obtained by approximating $|x|$ as $\sqrt{x^{2}+\varrho}-\sqrt{\varrho}$, where $\varrho$ is a positive constant. Hence, instead of (65) with (66), a smoothed constraint can be obtained,
$g({\mathbf{x}})$ $=\sum\nolimits_{k=1}^{n}\Big{(}\lambda\sqrt{[{\mathbf{x}}]^{2}_{k}+\varrho}-\lambda\sqrt{\varrho}-h_{\lambda,\theta}([{\mathbf{x}}]_{k})\Big{)}$ (67)
@@
where $g$ is smoothed version of $\bar{g}$ and $h$ is the smoothed version of of $\bar{h}$. Since the objective is already convex, its running surrogate ${\tilde{f}}$ can be constructed using ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ and (6) as
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ (68)
@@
where ${\boldsymbol{\xi}}_{t}=\{{\mathbf{a}}_{t},b_{t}\}$. On the other hand, the surrogate of $h$ can be constructed as,
$\tilde{h}_{\lambda,\theta}(x,x_{t})$ $=h_{\lambda,\theta}(x_{t})+\nabla h_{\lambda,\theta}(x_{t})(x-x_{t})$ (69)
@@
=$=\begin{cases}\frac{x_{t}x}{\theta}-\frac{x_{t}^{2}}{2\theta},&|x_{t}|\leq\theta\lambda\\ \frac{\lambda\varrho}{\sqrt{x_{t}^{2}+\varrho}}-\lambda\sqrt{\varrho}-\frac{\theta\lambda^{2}}{2}+\frac{\lambda x_{t}x}{\sqrt{x_{t}^{2}+\varrho}},&|x_{t}|>\theta\lambda.\\ \end{cases}$ (70)
Hence, the surrogate of the smoothed constraint (67) at point ${\mathbf{x}}_{t}$ becomes
${\tilde{g}}({\mathbf{x}},{\mathbf{x}}_{t})$ $=\sum\nolimits_{k=1}^{n}\Big{(}\lambda\sqrt{[{\mathbf{x}}]^{2}_{k}+\varrho}-\tilde{h}_{\lambda,\theta}([{\mathbf{x}}]_{k},[{\mathbf{x}}_{t}]_{k})-\lambda\sqrt{\varrho}+\tfrac{L}{2}([{\mathbf{x}}]_{k}-[{\mathbf{x}}_{t}]_{k})^{2}\Big{)}$ (71)
@@
It is worthy to note that 71 fulfills A9. We compare the performance of our algorithm with that of the LCPP method, which was proposed in [9] for solving stochastic non-convex optimization problems with sparsity inducing non-convex constraints. It is proximal point algorithm that solves the a sequence of convex subproblems with gradually relaxed constraint levels.
@@
MNIST $60000$ $10000$ $784$
gisette $6000$ $1000$ $5000$
For our simulations use two dfatasets: MNIST and gisette, details of which are summarized in Table II. For the MNIST dataset, we consider the binary classification problem of classifying the handwritten digit 5 from the other digits. As the dataset consists of 10 classes this can be achieved by assigning a label $+1$ to digit $5$ and $-1$ to other digits. Likewise, gisette dataset also consists handwritten digits, but here the classification is between the often confused digits 4 and 9.
For MCP we take $\lambda=2$ and $\theta=5$, which correspond to one of the parameter values used in [9]. The following CoSTA parameters were used after tuning them for best performance: (i) $\mu=0.05,{\bar{k}}=0.0051,c=6\times 10^{5},w=9000$ for gisette; and (ii) $\mu=0.06,{\bar{k}}=0.0018,c=1.4\times 10^{6},w=38000$ for MNIST.
@@
TABLE III: Classification accuracy of different algorithms for sparse logistic regression (where the entries are train accuracy/ test accuracy)
@@
MNIST train $90.5\pm 1.4$ $94\pm 0.56$
test $90.54\pm 1.51$ $94.1\pm 0.74$
gisette train $99.7\pm.05$ $99.44\pm 0.11$
test $97.36\pm 0.5$ $97.72\pm 0.32$
(a) $T_{f}=10$
(b) $T_{f}=15$
(c) $T_{f}=20$
Figure 4: Trajectory (top) and energy evolution vs iterations (bottom) plots for $T=30$, $\sigma=0.2$, environmental configuration 4, and different values of $T_{f}$.
TABLE IV: Average time per iteration in seconds for $\sigma$ = 0.2, $T_{f}$ = 15, environmental configuration 4, and different values of $T$.
@@ V-B Energy-Optimal Trajectory Planning
Continuing the problem discussed in Sec. II, we apply CoSTA and CSSCA [15] to solve the energy-efficient trajectory planning problem (1)-(3) in a simulated 2D ocean environment. The ocean currents at a point ${\mathbf{x}}=[x_{1}~{}x_{2}]^{\mathsf{T}}$ are simulated using the following equations
@@
We generate current predictions by adding unbiased noise as follows:
$\vartheta({\mathbf{x}},{\boldsymbol{\xi}})=\vartheta({\mathbf{x}})(\mathbf{I}+\text{diag}(\mathbf{e}({\boldsymbol{\xi}}))),$ (73)
where $\mathbf{e({\boldsymbol{\xi}})}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. As stated earlier, neither the true ocean currents nor the noise model is known to the vehicle, which can only query the predictor for noisy realizations $\vartheta({\mathbf{x}},{\boldsymbol{\xi}})$ of the currents.
We carry out Monte Carlo simulations on a Intel(R) Xeon(R) E3-1226 CPU running Ubuntu 20.04 LTS with 32GB RAM. We modeled the environment and implemented the algorithms in MATLAB 2024a [49]. Results are shown for multiple environments consisting of two agents and an obstacle. Each environment has a different starting, goal, and obstacle locations. Of these, the third environment, shown in Fig. 3(c) is particularly challenging, as the ocean currents tend to push one of the agents into the obstacle, requiring more precise navigation. We also vary the values of number of way points $T$, mission time $T_{f}$, and the noise parameter $\sigma$, so as to compare the performance of the proposed CoSTA algorithm with that of CSSCA across diverse settings.
The parameters of CSSCA ($\mu$, $\rho$, and $\gamma$ [15]) and CoSTA ($\mu$, ${\bar{k}}$, $c$, and $w$) were tuned for best possible performance after running the algorithms for initial 100 iterations. The results are averaged over 10 Monte Carlo runs for the following values of problem parameters: $\omega=0.8$, $r^{0}=0.7$m, $r=0.1$m, and $v_{i}^{\max}=1$m/s for $i\in\{1,2\}$. The initial trajectory for both the algorithms is generated by directly solving the non-convex feasibility problem (1b)-(1d),(3) using the MATLAB’s non-linear optimization function fmincon. The subproblems for CoSTA and CSSCA are Quadratic Constrained Quadratic Programs (QCQP) and solved them using MATLAB function fmincon [50] which uses the interior-point method to solve the convex subproblem.
@@
We also plot trajectories and evolution of energy cost vs iterations for different mission times $T_{f}$ in Fig. 4. When the mission times are larger, both the algorithms generate longer paths so as to leverage the additional time and obtain trajectories with lower energy. Finally, Table IV shows the per iteration time for CoSTA and CSSCA for different number of waypoints $T$. As the number of waypoints increases, the size of each subproblem increases and hence more time is required. Note that since the subproblems for CoSTA and CSSCA are similar, their per-iteration time is also similar. Hence, we can conclude that the reason for speedup of CoSTA is that it requires fewer iterations due to its use of momentum.

### v4

@@ IV Complexity analysis
Under milder MSS assumption A6 and using a non-adaptive step size, along with the bounds established in Corollary 2 and Lemma 4, there exists some $1\leq t\leq{\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ and ${\hat{{\boldsymbol{\lambda}}}}_{t}\in{\mathbb{R}}^{I}_{+}$, ${\hat{{\boldsymbol{\nu}}}}_{t}\in{\mathbb{R}}^{J}_{+}$ such that the point $({{\hat{{\mathbf{x}}}}_{t}},{\hat{{\boldsymbol{\lambda}}}}_{t},{\hat{{\boldsymbol{\nu}}}}_{t})$ is $\epsilon$-KKT optimal.
Under the adaptive step-size assumption, the SFO complexity bound in Theorem 2 matches that of STORM [20] and nearly matches Pstorm [33], STORM+ [30] achieving a complexity of $\mathcal{O}(\epsilon^{-3/2})$ for the unconstrained case. Notably, while STORM and Pstorm were proposed for unconstrained and convex-constrained problems, respectively, CoSTA extends this to non-convex constrained framework with a comparable SFO complexity under similar assumptions. Furthermore, under the non-adaptive step-size setting, CoSTA attains the same convergence rate but under milder MSS assumption. A key distinction of CoSTA, however, lies in its ability to achieve a near-optimal rate (i.e., ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$) under adaptive step sizes, a feature that distinguishes it from [14], which employs non-adaptive step sizes. It is remarked that the non-asymptotic results in Theorem 2 and Corollary 3 can also be used to infer the asymptotic convergence of the corresponding quantities.
@@ V-A Sparse Binary Classification
The problem of finding a sparse binary classifier can be written in the form of ($\mathcal{P^{\prime}}$) with the objective being the logistic loss function, i.e., $f({\mathbf{x}},\{{\mathbf{a}},b\})=\log(1+e^{-b{\mathbf{a}}^{\mathsf{T}}{\mathbf{x}}})$ and the constraint is given by $\bar{g}({\mathbf{x}})\leq\tau$, where $\bar{g}$ is a sparsity-inducing function such as the $\ell_{1}$ norm or the minimax concave penalty (MCP) [9]. Here, the random variable denotes independent observations $\{{\mathbf{a}},b\}$. In [9], the constraint function $\bar{g}$ is chosen to be the MCP, defined as: $\bar{g}({\mathbf{x}})=\lambda\|{\mathbf{x}}\|_{1}-\sum\nolimits_{k=1}^{n}h_{\lambda,\theta}([{\mathbf{x}}]_{k})$ where $[{\mathbf{x}}]_{k}$ denotes the $k$-th entry of ${\mathbf{x}}$ and $h_{\lambda,\theta}$ is defined as:
$h_{\lambda,\theta}(x)=\begin{cases}\frac{x^{2}}{2\theta},&|x|\leq\theta\lambda\\ \lambda|x|-\frac{\theta\lambda^{2}}{2},&|x|>\theta\lambda.\\ \end{cases}$ (64)
Since $\bar{g}$ in (64) is non-smooth, we consider an alternative smooth constraint given by,
$g({\mathbf{x}})$ $=\sum\nolimits_{k=1}^{n}\Big{(}h_{\lambda,\varrho}([{\mathbf{x}}]_{k})-h_{\lambda,\theta}([{\mathbf{x}}]_{k})\Big{)}\leq\tau$ (65)
where $\varrho,\theta,\lambda>0$ and $\varrho\ll\theta$. This results in a smooth, non-convex function $g$ that can be explicitly written as
$g(x)=\begin{cases}\frac{x^{2}}{2}\Big{(}\frac{1}{\varrho}-\frac{1}{\theta}\Big{)},&|x|\leq\varrho\lambda\\ \lambda|x|-\frac{\varrho\lambda^{2}}{2}-\frac{x^{2}}{2\theta},&\varrho\lambda<|x|\leq\theta\lambda\\ \frac{(\theta-\varrho)\lambda^{2}}{2},&|x|>\theta\lambda.\end{cases}$ (66)
@@
Since the objective is already convex, its running surrogate ${\tilde{f}}$ can be constructed using ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ and (6) as
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ (67)
@@
$+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$
@@
where ${\boldsymbol{\xi}}_{t}=\{{\mathbf{a}}_{t},b_{t}\}$. The constraint $g$ in (65) is non-convex and hence, the surrogate ${\tilde{g}}$ at point ${\mathbf{x}}_{t}$ can be constructed as,
${\tilde{g}}({\mathbf{x}},{\mathbf{x}}_{t})$ (68)
$=\sum\nolimits_{k=1}^{n}\Big{(}\tilde{g}_{lp}([{\mathbf{x}}]_{k},[{\mathbf{x}}_{t}]_{k})+\tfrac{L}{2}([{\mathbf{x}}]_{k}-[{\mathbf{x}}_{t}]_{k})^{2}\Big{)}\leq\tau$
where $\tilde{g}_{lp}$ is defined as
$\tilde{g}_{lp}(x,x_{t})=g(x_{t})+\langle\nabla g(x_{t}),x-x_{t}\rangle$
@@
=$=\begin{cases}\frac{x_{t}x}{\varrho}-\frac{x_{t}x}{\theta}-\frac{x_{t}^{2}}{2\varrho}+\frac{x_{t}^{2}}{2\theta},&|x_{t}|\leq\varrho\lambda\\ \lambda\cdot x\text{sign}(x_{t})-\frac{x_{t}x}{\theta}-\frac{\varrho\lambda^{2}}{2}+\frac{x_{t}^{2}}{2\theta},&\varrho\lambda<|x_{t}|\leq\theta\lambda\\ \frac{(\theta-\varrho)\lambda^{2}}{2},&|x_{t}|>\theta\lambda.\end{cases}$ (69)
@@
Observe that 68 satisfies A9. We compare the performance of our algorithm with that of the LCPP method, which was proposed in [9] for solving stochastic non-convex optimization problems with sparsity inducing non-convex constraints. It is proximal point algorithm that solves the a sequence of convex subproblems with gradually relaxed constraint levels.
@@
MNIST 600006000060000 100001000010000 784784784
gisette 600060006000 100010001000 500050005000
For our simulations use two dfatasets: MNIST and gisette, details of which are summarized in Table II. For the MNIST dataset, we consider the binary classification problem of classifying the handwritten digit 5 from the other digits. As the dataset consists of 10 classes this can be achieved by assigning a label +1+1+ 1 to digit 555 and 1-1- 1 to other digits. Likewise, gisette dataset also consists handwritten digits, but here the classification is between the often confused digits 4 and 9.
For MCP we take $\lambda=2$ and $\theta=5$, which correspond to one of the parameter values used in [9]. The following CoSTA parameters were used after tuning them for best performance: (i) $\mu=0.05,{\bar{k}}=0.0051,c=6\times 10^{5},w=9000,\tau=0.05n$ for gisette; and (ii) $\mu=0.06,{\bar{k}}=0.0018,c=1.4\times 10^{6},w=38000,\tau=0.1n$ for MNIST.
@@
TABLE III: Classification accuracy of different algorithms for sparse logistic regression (where the entries are train accuracy/ test accuracy).
@@
MNIST train 90.5±1.490.5$\pm$ 1.490.5 ± 1.4 94±0.5694$\pm$ 0.5694 ± 0.56
test 90.54±1.5190.54$\pm$ 1.5190.54 ± 1.51 94.1±0.7494.1$\pm$ 0.7494.1 ± 0.74
gisette train 99.7±.0599.7$\pm$.0599.7 ± .05 99.44±0.1199.44$\pm$ 0.1199.44 ± 0.11
test 97.36±0.597.36$\pm$ 0.597.36 ± 0.5 97.72±0.3297.72$\pm$ 0.3297.72 ± 0.32
TABLE IV: Average time per iteration in seconds for $\sigma$ = 0.2, $T_{f}$ = 15, environmental configuration 4, and different values of $T$ for Case S1.
@@ V-B Energy-Optimal Trajectory Planning
Continuing the problem discussed in Sec. II, we benchmark our proposed algorithm CoSTA against classical optimization based approach in CSSCA [15] and sampling based methods STOMP [49] and EESTO [50] on the energy-efficient trajectory planning problem (1)-(3) in a simulated 2D ocean environment. Monte Carlo simulations were performed on an Intel(R) Xeon(R) E3-1226 CPU running Ubuntu 20.04 LTS with 32GB RAM. We modeled the environment and implemented the algorithms in MATLAB 2024a [51].
We consider two cases of simulated ocean current setting with different complexity. For case S1 we consider a simple ocean current field with a single obstacle. The second case S2 considers navigation through a narrow passage with a complex ocean current field and multiple obstacles. For each case, the base ocean current field $\vartheta({\mathbf{x}})$ is generated differently. From the base ocean currents we generate current ensemble predictions $\vartheta({\mathbf{x}},{\boldsymbol{\xi}})$ by adding noise as follows:
$\vartheta({\mathbf{q}},{\boldsymbol{\xi}})=\vartheta({\mathbf{q}})(\mathbf{I}+\text{diag}(\mathbf{e}({\boldsymbol{\xi}}))),$ (70)
where $\mathbf{e({\boldsymbol{\xi}})}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. As stated earlier, neither the true ocean currents nor the noise model is known to the vehicle, which can only query the predictor for noisy realizations $\vartheta({\mathbf{q}},{\boldsymbol{\xi}})$ of the currents. We discuss the setting and results for both the cases in detail in the following subsections.
Case S1 - Convergence Benchmark against optimization based algorithms
Here we consider a simple setting with one obstacle where the ocean currents at a point ${\mathbf{x}}=[x_{1}~x_{2}]^{\mathsf{T}}$ are simulated using the following:
@@
For S1 we compare the performance of proposed algorithm CoSTA with that of CSSCA across diverse settings. In this case, we run simulations for multiple environments consisting of two agents and an obstacle. Each environment has a different starting, goal, and obstacle locations. Of these, the third environment, shown in Fig. 3(c) is particularly challenging, as the ocean currents tend to push one of the agents into the obstacle, requiring more precise navigation. We also vary the values of number of way points $T$, mission time $T_{f}$, and the noise parameter $\sigma$.
The parameters of CSSCA ($\mu$, $\rho$, and $\gamma$ [15]) and CoSTA ($\mu$, ${\bar{k}}$, $c$, and $w$) were tuned for best possible performance after running the algorithms for initial 100 iterations. The results are averaged over 10 Monte Carlo runs for the following values of problem parameters: $\omega=0.8$, $r^{0}=0.7$m, $r=0.1$m, and $v^{\max}=1$m/s. The initial trajectory for both the algorithms is generated by directly solving the non-convex feasibility problem (1b)-(1d),(3) using the MATLAB’s non-linear optimization function fmincon. The subproblems for CoSTA and CSSCA are Quadratic Constrained Quadratic Programs (QCQP) and solved using MATLAB function fmincon [52] which uses the interior-point method to solve the convex subproblem.
@@
Table IV shows the per iteration time for CoSTA and CSSCA for different number of waypoints $T$. As the number of waypoints increases, the size of each subproblem increases and hence more time is required. Note that since the subproblems for CoSTA and CSSCA are similar, their per-iteration time is also similar. Hence, we can conclude that the reason for speedup of CoSTA is that it requires fewer iterations due to its use of momentum.
Case S2 - Effectiveness Benchmark against planning algorithms
For this case, we consider a more challenging scenario of narrow passage navigation through multiple obstaces. Currents are simulated using viscous lamb vortices [53] which are generated using superposition of one point vortex solutions. The model with $M$ vortices is mathematically described by :
$\vartheta({\mathbf{x}})=\sum_{m=1}^{M}\dfrac{\Omega_{m}({\mathbf{x}}-{\mathbf{q}}_{m})}{2\pi\left\|{\mathbf{x}}-{\mathbf{q}}_{m}\right\|^{2}}\left(1-\exp\left(-\dfrac{\left\|{\mathbf{x}}-{\mathbf{q}}_{m}\right\|^{2}}{\delta_{m}^{2}}\right)\right)$ (72)
with $\Omega_{m}=\begin{bmatrix}0&-\omega_{m}\\ \omega_{m}&0\end{bmatrix}$, where ${\mathbf{q}}_{m}$ is the centre of the $m$-th vortex and $\omega_{m}$, $\delta_{m}$ are parameters related to its strength and radius, respectively. In a 200m $\times$ 200m environment, we consider ocean model with $M=5$ Lamb vortices, $\omega\in\{-20,-10,20\}$, and $\delta\in\{10,20\}$. Two agents with $r=5$m and $v^{\max}=1$m/s seek to navigate to their respective goal locations in $T_{f}=300s$ avoiding each other and 4 obstacles with radii $r^{o}\in\{10,20,30,35\}$ m.
The initial trajectory for CoSTA for this case is generated in a similar way to case S1 by directly solving the non-convex feasibility problem (1b)-(1d),(3) but this time with an energy-informed-AStar [^2] The energy-informed-AStar is simple AStar search with energy as travel cost[54] initialization. This results in a overall feasible initial trajectory for CoSTA.
TABLE V: Comparison for Case S2
Algorithm Energy (J) $\mathbb{E}$[Constraint Violation]
Initial guess 243.71$\pm$1.53 0
CoSTA 139.33$\pm$4.20 0
EESTO [50] 604.84$\pm$22.02 6.94
STOMP [49] 622$\pm$32.53 5.89
We benchmark CoSTA against sampling-based planning methods STOMP and EESTO. All the algorithms were independently tuned for best possible performance. For this case all algorithms are again run for 10 Monte Carlo runs. A sample trajectory obtained from one of the runs is shown in Fig. 4. Observe, in Fig. 4, for one of the agents CoSTA trajectory takes a long path along the flow to conserve energy whereas STOMP and EESTO get stuck in local minimum taking a shorter but flow opposing energy hungry trajectory. Sampling-based approaches such as STOMP and EESTO operate using average ocean current fields, as they are not equipped to incorporate ensemble current forecasts. Consequently, their performance tends to be suboptimal in this setting.
Table V reports the average energy consumption and average constraint violations over ocean ensembles for the competing algorithms. All algorithms produce collision-free trajectories; however, CoSTA achieves the lowest average energy consumption without any constraint violations, thereby demonstrating superior performance.
Figure 4: Trajectory plots for case S2 environment with 5 Lamb vortices.